# Longitudinal efficient adjustment sets for time-varying treatment effect estimation in nonparametric causal graphical models

### v1

@@ 5 Conclusion
In this article, we built on the work of Rotnitzky and Smucler [5] and proposed an alternative definition of sufficient time-dependent adjustment sets that takes into account potential simplifications to the identification formula using conditional independencies that can be read from the causal graph. We proposed two lemmas and a theorem that allow comparing the asymptotic variance of efficient estimators based on our definition of a sufficient time-dependent adjustment sets and showed that further variance reduction can be obtained as compared to estimators based on previous results. We provided two numerical illustrations of our results. In one of those examples, the optimal adjustment set varies according to the data-generating equations for the same DAG when employing Rotnitzky and Smucler’s criterion. In contrast, among the sufficient time-dependent adjustment sets that meet our definition, the optimal set seems to remain the same regardless of the data-generating equations for the same DAG. Based on those results, we conjecture that it is possible to identify an optimal time-dependent adjustment set based on a directed acyclic graph alone when considering our definition of a sufficient time-dependent adjustment set. A formal demonstration of this result is left for future work.
@@
Our current contribution has several important implications. For example, our results can be used directly by data-analysts when estimating the effect of a time-varying treatment in order to improve the precision of their estimates. Moreover, our results could be used in developing data-driven variable selection procedures. Indeed, several confounder selection approaches for a time-fixed covariate have been developed for targeting the optimal adjustment set (e.g.,[14, 15]). Extensions to the time-varying treatment setting may be possible based on the results of Rotnitzky and Smucler [5] and ours.
Despite the strengths of work, it is important to highlight some of its limitations. A first limitation is that we were not able to demonstrate (yet) our conjecture that our alternative definition of a sufficient time-dependent adjustment set leads to the possibility of uniquely identifying an optimal time-dependent adjustment set based on the DAG alone. A second limitation is that we have only considered DAGs where all variables are observed. In situations where some DAG vertices are unobservable, but some adjustment sets are observable, Henckel et al [4] and Rotnitzky and Smucler [5] have provided theoretical results indicating that it is not always possible to find a uniformly optimal set among observable adjustment sets.
@@
Our current work could be extended in several directions. In studies with time-dependent confounding, it is very common for data from some individuals to be missing or incomplete due to loss to follow-up or other factors (i.e., censored data). Ignoring these losses to follow-up when identifying the adjustment set may induce selection bias. Correa et al [16] proposed a necessary and sufficient graphical criterion to estimate causal effect estimation under confounding and selection bias for the cross-sectional case. Future work could investigate whether the results of Correa et al [16] can be adapted to the longitudinal case, in particular by considering the approach presented in this paper. Finally, to facilitate the use of our proposed rules, it would be greatly valuable to develop efficient algorithms and implement them in open-source software, similar to how DAGitty can be used to identify a minimal sufficient adjustment set based on a causal graph [17], or how the optAdjSet function of the pcalg package in R can be used to identify an optimal adjustment set based on a causal graph in the single treatment time point setting.

### v2

@@ 5 Conclusion
In this article, we built on the work of Rotnitzky and Smucler [5] and proposed an alternative definition of sufficient time-dependent adjustment sets that takes into account potential simplifications to the identification formula using conditional independencies that can be read from the causal graph. We proposed two lemmas and a theorem that allow comparing the asymptotic variance of efficient estimators based on our definition of sufficient time-dependent adjustment sets and showed that further variance reduction can be obtained as compared to estimators based on previous results. We have further demonstrated that it is possible to identify an optimal sufficient time-dependent adjustment set based on the graph alone when considering our definition. We provided two numerical illustrations of our results. In one of those examples, the optimal adjustment set varies according to the data-generating equations for the same DAG when employing Rotnitzky and Smucler’s definition [5]. In contrast, among the sufficient time-dependent adjustment sets that meet our definition, the optimal set remains the same regardless of the data-generating equations for the same DAG. The reduction in the variability of estimates allowed by our results was observed to be appreciable: the standard deviation of the estimates obtained using our optimal adjustment sets was between 10% and 20% lower than the standard deviation using the optimal sets according to previous results.
@@
Our current contribution has several important implications. For example, our results can be used directly by data-analysts when estimating the effect of a time-dependent treatment in order to improve the precision of their estimates. Moreover, our results can be used for developing data-driven variable selection procedures. Indeed, several confounder selection approaches for a time-fixed covariate have been developed for targeting the optimal adjustment set [13, 14]. Extensions to the time-varying treatment setting based on the results of Rotnitzky and Smucler [5] and ours are already underway [15].
@@
Our current work could be extended in several directions. First, we have only considered DAGs where all variables are observed. In situations where some DAG vertices are unobservable, but some adjustment sets are observable, Henckel et al [4] and Rotnitzky and Smucler [5] have provided theoretical results indicating that it is not always possible to find a uniformly optimal set among observable adjustment sets. It would interesting to similarly extend our results to situations where not all variables are observed. In studies with time-dependent confounding, it is also very common for data from some individuals to be missing or incomplete due to loss to follow-up or other factors (i.e., censored data). Ignoring these losses to follow-up when identifying the adjustment set may induce selection bias. Correa et al [16] proposed a necessary and sufficient graphical criterion to estimate causal effects under confounding and selection bias for the cross-sectional case. Our results directly extend to the case where the interest lies specifically in estimating the effect of a treatment regime under no censoring, by simply treating censoring nodes as treatment nodes. Finally, to facilitate the use of our proposed rules, it would be greatly valuable to develop efficient algorithms and implement them in open-source software, similar to how DAGitty can be used to identify a minimal sufficient adjustment set based on a causal graph [17], or how the optAdjSet function of the pcalg package in R can be used to identify an optimal adjustment set based on a causal graph in the single treatment time point setting.
Acknowledgements
The authors would like to thank Geneviève Lefebvre for her comments on earlier version of this manuscript. This work was supported by a Pierre Jacob Durand – Citoyens du monde scholarship from Université Laval to DA; a Discovery Grant from the Natural Sciences and Engineering Research Council of Canada to DT (RGPIN-2023-04911); Fonds de recherche du Québec – Santé research career awards to DT (https://doi.org/10.69777/312198) and JRG (https://doi.org/10.69777/330678); and a Canada Research Chair to MES.